# Signed Networks: theory, methods, and applications

### v1

@@
Fernando Diaz-Diaz Universidad Carlos III de Madrid, Departamento de Matemáticas, Grupo Interdisciplinar de Sistemas Complejos, 28911 Leganés, Spain Institute of Cross-Disciplinary Physics and Complex Systems, IFISC (UIB-CSIC), 07122 Palma de Mallorca, Spain Elena Candellone Department of Methodology and Statistics, Utrecht University, Utrecht, Netherlands Centre for Complex Systems Studies, Utrecht University, Utrecht, Netherlands Miguel A. González-Casado Universidad Carlos III de Madrid, Departamento de Matemáticas, Grupo Interdisciplinar de Sistemas Complejos, 28911 Leganés, Spain Emma Fraxanet Department of Engineering, Universitat Pompeu Fabra, Barcelona 08018, Spain Antoine Vendeville Sciences Po médialab, 75007 Paris, France Complex Systems Institute of Paris Ile-de-France CNRS, 75013 Paris, France Learning Planet Institute, Learning Transitions unit, CY Cergy Paris University, 75004 Paris, France Irene Ferri Departament de Física de la Matèria Condensada, Universitat de Barcelona (UB), c. Martí i Franquès, 1, 08028 Barcelona, Spain Institut de Recerca en Sistemes Complexos (UBICS), Universitat de Barcelona (UB), Barcelona, Spain The Roux Institute, Network Science Institute, Northeastern University, 04101 Portland, ME USA Andreia Sofia Teixeira BRAN Lab, Network Science Institute, Northeastern University London, London E1W 1LP, United Kingdom LASIGE, Departamento de Informática, Universidade de Lisboa, Campo Grande 1749-016, Lisboa, Portugal
@@ 1 Introduction
Signed networks represent systems of interconnected entities, in which connections can be positive or negative. They offer a powerful framework for studying systems in which interactions are not merely present or absent, but qualitatively different: cooperative or antagonistic, supportive or conflicting, excitatory or inhibitory. Traditional network approaches that ignore negative ties or treat them as absent connections miss essential dynamics: how antagonism drives faction formation, how the interplay of cooperation and conflict creates polarisation, and how systems organise around the tension between positive and negative relationships. Individual disciplines have long recognised the importance of antagonistic relationships (see Figure 1). Ecologists study competition and predation alongside mutualism [allesinaStabilityCriteriaComplex2012]. Political scientists examine rivalries and conflicts as central features of international systems [hafner-burtonNetworkAnalysisInternational2009]. Sociologists have studied social systems in which positive edges can represent friendship, trust, or alliance, whereas negative edges may encode hostility, distrust, or rivalry [ruiz-garciaTriadicInfluenceProxy2023]. In economics, researchers have studied financial volatility and how positive correlations between assets, indicating shared exposure to a certain risk factor, and negative correlations signal diversification or hedging opportunities [souto2018capturing, bartesaghi2025global]. In Psychology and Cognitive Science, sentiments, beliefs, and attitudes are studied through networks, where positive edges represent mutually reinforcing or consistent beliefs, and negative edges may represent conflicting or opposing beliefs [teixeira2021revealing, aiyappaEmergenceSimpleComplex2024]. What has been missing is a unified vocabulary and framework that allows a shared formalism to address these distinct fields. Signed networks provide this framework. By representing systems as graphs where edges carry explicit positive or negative signs, we can identify structural patterns and dynamical processes that go beyond a single application. A signed network of international alliances and rivalries, as well as a social network of friendship and enmity, may represent different systems, but they can exhibit the same mathematical properties, be analysed with the same computational methods, and follow similar organising principles.
@@
The intellectual foundations of signed networks emerged over 75 years ago from the concept of balance. Psychologist Fritz Heider [heiderAttitudesCognitiveOrganization1946] proposed that individuals experience psychological tension in certain configurations of relationships. For instance, when a person has two friends who are enemies with each other, or likes someone who dislikes their friend. Social psychologist Dorwin Cartwright and mathematician Frank Harary [cartwrightStructuralBalanceGeneralization1956] formalised this insight: a network is “balanced” when every cycle contains an even number of negative relationships. This simple principle has profound implications. Balanced networks naturally organise into mutually antagonistic factions, with positive edges within groups and negative edges between them. This fundamental insight remains central to signed network research today. For several decades, however, empirical work remained limited to small-scale sociometric studies [sampsonNovitiatePeriodChange1968, newcombAcquaintanceProcess1961], constrained by the difficulty of collecting signed relational data and limited computational tools. The field transformed in the early 2000s with the rise of modern network science as a discipline and with large-scale digital platforms that explicitly recorded signed interactions, such as trust/distrust on Epinions, friend/foe tags on Slashdot, or support/oppose votes on Wikipedia [leskovecPredictingPositiveNegative2010]. These datasets enabled researchers to test classical theories at scale and develop new computational methods. The 2010s brought rapid methodological innovation: graph neural networks adapted for signed edges[derrSignedGraphConvolutional2018], spectral methods exploiting signed Laplacian properties [kunegisSpectralAnalysisSigned2010], stochastic block models for statistical inference [jiangStochasticBlockModel2015], maximum entropy frameworks for null models [galloTestingStructuralBalance2024], and walk-based balance measures [estradaWalkbasedMeasureBalance2014, talagaPolarizationMultiscaleStructural2023]. Interest spread across disciplines, such as physics, neuroscience, ecology, computer science, and political science, each recognising that antagonistic relationships fundamentally shape system organisation. Today, signed network science research is more active than ever, with hundreds of papers published annually across diverse fields, powerful computational methods for analysis and prediction, and growing availability of large-scale signed datasets. Applications range from understanding online polarisation [vendevilleModelingEchoChamber2025, fraxanetUnpackingPolarizationAntagonism2024], international conflicts [diaz-diazMathematicalModelingLocal2024], to the analysis of brain connectivity [tanner2024multi]. This growth reflects both the field’s maturity and its continued relevance to pressing questions about how cooperation and conflict shape complex systems. Yet, this rapid, distributed growth has created its own challenges.
@@
This book provides a comprehensive, systematic treatment of signed networks that consolidates foundational theories, clarifies methodological choices, identifies persistent misconceptions, and discusses paths forward. We treat signed networks as a unified analytical framework while respecting that different domains may require different approaches. We emphasise methodological rigour and transparency, discussing assumptions, limitations, and appropriate contexts rather than simply cataloguing methods. We integrate across levels of analysis—static structure and temporal dynamics, processes on and of networks, microscopic mechanisms and macroscopic patterns. We are selective rather than exhaustive, focusing on foundational concepts, broadly useful methods, and findings that reveal general principles. The book builds systematically. Chapter 2 establishes mathematical foundations: formalising signed networks and presenting the adaptation of classical metrics. Chapter 3 addresses balance theory, distinguishing cognitive from structural balance and presenting measures for quantifying balance. Chapter 4 explores structural inference: null models, embeddings, and algorithms for sign prediction and community detection. Chapter 5 examines dynamics: opinion formation, network evolution, and co-evolutionary frameworks. Chapter 6 tackles data collection and surveys available datasets. Chapter 7 addresses common misconceptions and errors. In Chapter 8, we discuss open challenges and promising directions. Our aim with this book is to make signed network science more accessible without sacrificing rigour, more unified without erasing important distinctions, and more self-aware about its limitations, helping researchers effectively use this powerful lens for understanding how positive and negative relationships jointly shape complex systems.
Figure 1: Examples of signed networks across different domains. In social systems, nodes represent individuals and edges denote friendship/enmity or trust/distrust relationships. In ecological systems, edges capture competitive, mutualistic, or predator–prey dynamics. In economic systems, assets are connected based on correlations between their price time series. In psychological systems, nodes represent beliefs or attitudes, and signed edges encode whether they reinforce or contradict each other. In biological systems, interactions between genes and proteins can be represented through activation/inhibition links. In neuroscience, signed networks can represent either excitatory/inhibitory synapses between individual neurons or functional relationships between brain regions. Finally, in international relations, countries may be connected by alliances or conflicts. Figures adapted from [dorresteijnIncorporatingAnthropogenicEffects2015, aiyappaEmergenceSimpleComplex2024, markuTimeseriesTranscriptomicsGene2023, whitfield-gabrieliDefaultModeNetwork2012, diaz-diazMathematicalModelingLocal2024].
@@ 2.1 Definition and representation of signed networks
Let us introduce the fundamental mathematical concepts and terms, and some of the key concepts used throughout this book. From a mathematical point of view, networks (whether signed or unsigned) are modelled as graphs [newmanNetworks2018]. A graph $G=(V,E)$ consists of a set $V=\{v_{i}\}$ of nodes (also called vertices) and a set $E=\{e_{ij}\}$ of edges (also called links) that connect pairs of nodes. An edge connecting the nodes $v_{i}$ and $v_{j}$ is denoted by $e_{ij}=(v_{i},v_{j})$. The number of nodes is denoted by $n$ and the number of edges by $m$. Nodes can represent entities such as people, genes, countries, or more abstract constructs like stocks or time series. Edges capture pairwise interactions between these nodes: social connections between individuals, regulatory relationships between genes, or correlations between time series, among others.
On a signed graph, edges carry a sign attribute [zaslavskySignedGraphs1982] expressing a negative or positive type of relation between the nodes. Beyond signs, edges may also have directionality to allow for the modelling of asymmetrical relations between nodes [mackay2020directed], and weights to quantify varying levels of importance across relations [barratArchitectureComplexWeighted2004, newmanAnalysisWeightedNetworks2004]. The interplay between these attributes can lead to rich emergent phenomena; for instance, biological circuits leverage both sign and directionality to generate oscillatory dynamics [alon2019introduction]. In this book, we focus on the specific role of signed edges, and we avoid details regarding the role of directionality and weights.
@@
The adjacency matrix enables the application of techniques from linear algebra and spectral graph theory to analyse graphs [biggsAlgebraicGraphTheory1993, zaslavskyMatricesTheorySigned2013]. When the graph is unsigned, the $(i,j)$th element of $A^{k}$ counts the number of walks of length $k$ between nodes $v_{i}$ and $v_{j}$. When the graph is signed, it counts instead the difference between the number of positive and negative walks. These are given respectively by
@@ 2.2 Basic metrics
Centrality measures allow us to characterise local and global properties of networks. However, typical network metrics, such as degree, centrality measures, shortest-path distances, clustering coefficients, and network motifs [newmanNetworks2018], are not straightforward to generalise for signed networks. Some of these cease to be well-defined (e.g. the shortest-path distance), while others admit different generalisations suitable to different contexts. In this section, we explore how to approach each of these metrics and centrality measures in signed networks in a principled manner.
@@ 2.2.1 Signed degree and its distribution
In an unsigned network, the degree of a node $v_{i}$, denoted by $k_{i}$, counts the number of edges adjacent to the node. In a social network, the degree usually quantifies the popularity or influence of a person. Studying the distribution of degrees $P(k)$ can then help answer questions such as whether social influence is concentrated in the hands of a few or more equally spread in the population. When the network is signed, the degree is not characterised by a single number, but by a pair $(k^{+},k^{-})$, denoting the number of positive and negative edges incident to the node, respectively. The two-dimensional nature of the degree causes the degree distribution of a signed network to become bivariate $P(k^{+},k^{-})$. Analysing or visualising bivariate distributions can be challenging, so it is common to consider the marginal distributions, $P(k^{+})$ and $P(k^{-})$. However, one should keep in mind that marginalising erases any correlation between $k^{+}$ and $k^{-}$, which may contain valuable structural information. An alternative is to analyse the absolute or underlying degree $k=k^{+}+k^{-}$, and the net or signed degree $k_{\text{net}}=k^{+}-k^{-}$, through their joint probability distribution $P(k,k_{\text{net}})$ and its marginals. Asymmetries have been found between the distributions $P(k^{+})$ and $P(k^{-})$ [szellMeasuringSocialDynamics2010], suggesting fundamental differences in the formation mechanisms of positive and negative links, and challenging the use of $P(k)$ as an appropriate coarse-graining.
@@ 2.2.2 Signed clustering coefficient
The (unsigned) clustering coefficient measures local cohesiveness by quantifying the abundance of triangles. This is a key structural property of networks that plays a central role in the famous small-world phenomenon [Watts1998]. For the signed case, the (global) clustering coefficient can be computed as [kunegisSlashdotZooMining2009]:
@@
where $k_{i}$ is the absolute degree of node $v_{i}$. This index captures the difference between the number of positive triangles $t^{+}$ and negative triangles $t^{-}$, with the remaining terms [^1] In the presence of self-loops, one must use the normalisation constant $k_{i}^{2}/2$ instead [kunegisSlashdotZooMining2009]. ensuring that it remains bounded between +1+1 and 1-1. Clustering can also be studied at the node level, quantifying the difference between the number of positive triangles ($t^{+}$) and negative triangles ($t^{-}$) containing node $v_{i}$, relative to all two-paths passing through $v_{i}$ [costantiniGeneralizationClusteringCoefficients2014]:
@@
Both $C$ and $C_{i}$ penalise negative triangles ($t^{-}$) and unclosed triangles. To isolate the effect of negative triangles while controlling for network topology, Kunegis introduced the relative clustering coefficient [kunegisSlashdotZooMining2009], defined as the ratio between the signed and unsigned clustering coefficients:
@@ 2.2.3 Signed distances
Another fundamental concept in network analysis is the distance between pairs of nodes. This measure plays a central role in various applications, including the definition of various centrality metrics [newmanNetworks2018, borgattiCentralityNetworkFlow2005], community detection [newmanFindingEvaluatingCommunity2004], and network geometry [bogunaNetworkGeometry2021]. Formally, a distance is a function that takes two nodes as input and returns a real number that satisfies four key properties.
@@
In the case of signed networks, however, these measures fail to satisfy the properties of a distance [diazdiazthesis]. To understand the reasons, consider a graph made of three nodes $v_{1},v_{2},v_{3}$, the first two linked by a negative edge and the last two by a positive one. Using the minimum-weight distance we obtain $d(v_{1},v_{2})<0$, violating the non-negativity condition, and $d(v_{1},v_{3})=0$, contradicting the identity of indiscernibles. These issues are not just theoretical; they can result in algorithms selecting infinitely long paths as optimal routes.
A natural way to overcome these limitations is to use generalised distances [bogunaNetworkGeometry2021]. These distances extend beyond simple shortest-path measures by incorporating the collective influence of all paths between two nodes, often through dynamic processes such as diffusion, random walks, or epidemic spreading. One prominent example is the resistance distance, which originated from electrical circuit theory [chandraElectricalResistanceGraph1989]. Here, the distance between two nodes corresponds to the effective resistance in a hypothetical circuit where a unit current (1 amp) is injected into $v_{i}$ and extracted from $v_{j}$, with every edge acting as a 1-ohm resistor. Alternatively, this measure has a probabilistic interpretation: it is equal to the expected commuting time for a random walker travelling from $v_{i}$ to $v_{j}$ and back. Mathematically, the resistance distance is given by
@@
where $L$ is the graph Laplacian [devriendtEffectiveResistanceMore2022], discussed in detail in Section 2.3, and $L^{\dagger}$ is its Moore-Penrose pseudoinverse, defined as: $L^{\dagger}=(L+\gamma J)^{-1}-\gamma^{-1}J$, with $\gamma$ a regularization parameter and $J$ an all-ones matrix [^3] The Moore–Penrose pseudoinverse generalizes the notion of the matrix inverse and remains well defined even when some eigenvalues are equal to zero, as in the Laplacian. For nonsingular matrices, it coincides with the ordinary inverse.. The function $d_{res}$ satisfies all properties of a distance, even when including negative edges [kunegisApplicationsStructuralBalance2014]. Moreover, it is Euclidean, meaning it can be interpreted as the usual “straight-line length” in some abstract Euclidean space.
The effective resistance is not the only generalised distance that retains good properties in signed networks. Other relevant ones include the diffusion distance [dedomenicoDiffusionGeometryUnravels2017], which captures how similar two nodes are based on the spread of a diffusive process over time; and the communicability distance [estradaComplexNetworksEuclidean2012, diazdiaz2025], which measures the ease of information transfer between nodes using the exponential of the adjacency matrix.
Despite the mathematical challenges discussed, algorithms exist to compute the minimum-weight distance in signed directed graphs. Two important examples are the Bellman-Ford and Floyd-Warshall algorithms [cormenIntroductionAlgorithmsFourth2022]. Both algorithms are based on dynamic programming and are preferred over Dijkstra’s algorithm when negative edges are present. Apart from computing minimum-weight paths, these algorithms also detect $\Sigma$-negative cycles (cycles where the sum of edge weights is negative) [^4] Note that this notion is different from negative cycles, where the product of edge weights is negative. $\Sigma$-negative cycles appear in diverse applications. For instance, in exchange-rate networks, currencies can be represented as nodes and exchange rates as edge weights (typically using the negative logarithm of the rate). In this setting, a $\Sigma$-negative cycle indicates an arbitrage opportunity..

### v2

@@
Fernando Diaz-Diaz Universidad Carlos III de Madrid, Departamento de Matemáticas, Grupo Interdisciplinar de Sistemas Complejos, 28911 Leganés, Spain Institute of Cross-Disciplinary Physics and Complex Systems, IFISC (UIB-CSIC), 07122 Palma de Mallorca, Spain Elena Candellone Department of Methodology and Statistics, Utrecht University, Utrecht, Netherlands Centre for Complex Systems Studies, Utrecht University, Utrecht, Netherlands Miguel A. González-Casado Universidad Carlos III de Madrid, Departamento de Matemáticas, Grupo Interdisciplinar de Sistemas Complejos, 28911 Leganés, Spain Emma Fraxanet Department of Engineering, Universitat Pompeu Fabra, Barcelona, Spain Computational Social Sciences and Humanities Lab, Barcelona Supercomputing Center, Barcelona, Spain Antoine Vendeville Sciences Po médialab, 75007 Paris, France Complex Systems Institute of Paris Ile-de-France CNRS, 75013 Paris, France Irene Ferri Departament de Física de la Matèria Condensada, Universitat de Barcelona (UB), c. Martí i Franquès, 1, 08028 Barcelona, Spain Institut de Recerca en Sistemes Complexos (UBICS), Universitat de Barcelona (UB), Barcelona, Spain The Roux Institute, Network Science Institute, Northeastern University, 04101 Portland, ME USA Andreia Sofia Teixeira BRAN Lab, Network Science Institute, Northeastern University London, London E1W 1LP, United Kingdom LASIGE, Departamento de Informática, Universidade de Lisboa, Campo Grande 1749-016, Lisboa, Portugal
@@ 1 Introduction
Signed networks represent systems of interconnected entities, in which connections can be positive or negative. They offer a powerful framework for studying systems in which interactions are not merely present or absent, but qualitatively different: cooperative or antagonistic, supportive or conflicting, excitatory or inhibitory. Traditional network approaches that ignore negative ties or treat them as absent connections miss essential dynamics: how antagonism drives faction formation, how the interplay of cooperation and conflict creates polarisation, and how systems organise around the tension between positive and negative relationships. Individual disciplines have long recognised the importance of antagonistic relationships (see Figure 1). Ecologists study competition and predation alongside mutualism [7]. Political scientists examine rivalries and conflicts as central features of international systems [160]. Sociologists have studied social systems in which positive edges can represent friendship, trust, or alliance, whereas negative edges may encode hostility, distrust, or rivalry [332]. In economics, researchers have studied financial volatility and how positive correlations between assets, indicating shared exposure to a certain risk factor, and negative correlations signal diversification or hedging opportunities [361, 33]. In Psychology and Cognitive Science, sentiments, beliefs, and attitudes are studied through networks, where positive edges represent mutually reinforcing or consistent beliefs, and negative edges may represent conflicting or opposing beliefs [381, 4]. What has been missing is a unified vocabulary and framework that allows a shared formalism to address these distinct fields. Signed networks provide this framework. By representing systems as graphs where edges carry explicit positive or negative signs, we can identify structural patterns and dynamical processes that go beyond a single application. A signed network of international alliances and rivalries, as well as a social network of friendship and enmity, may represent different systems, but they can exhibit the same mathematical properties, be analysed with the same computational methods, and follow similar organising principles.
@@
The intellectual foundations of signed networks emerged over 75 years ago from the concept of balance. Psychologist Fritz Heider [177] proposed that individuals experience psychological tension in certain configurations of relationships. For instance, when a person has two friends who are enemies with each other, or likes someone who dislikes their friend. Social psychologist Dorwin Cartwright and mathematician Frank Harary [55] formalised this insight: a network is “balanced” when every cycle contains an even number of negative relationships. This simple principle has profound implications. Balanced networks naturally organise into mutually antagonistic factions, with positive edges within groups and negative edges between them. This fundamental insight remains central to signed network research today. For several decades, however, empirical work remained limited to small-scale sociometric studies [339, 292], constrained by the difficulty of collecting signed relational data and limited computational tools. The field transformed in the early 2000s with the rise of modern network science as a discipline and with large-scale digital platforms that explicitly recorded signed interactions, such as trust/distrust on Epinions, friend/foe tags on Slashdot, or support/oppose votes on Wikipedia [231]. These datasets enabled researchers to test classical theories at scale and develop new computational methods. The 2010s brought rapid methodological innovation: graph neural networks adapted for signed edges[86], spectral methods exploiting signed Laplacian properties [224], stochastic block models for statistical inference [202], maximum entropy frameworks for null models [138], and walk-based balance measures [106, 374]. Interest spread across disciplines, such as physics, neuroscience, ecology, computer science, and political science, each recognising that antagonistic relationships fundamentally shape system organisation. Today, signed network science research is more active than ever, with hundreds of papers published annually across diverse fields, powerful computational methods for analysis and prediction, and growing availability of large-scale signed datasets. Applications range from understanding online polarisation [391, 131], international conflicts [90], to the analysis of brain connectivity [380, 89]. This growth reflects both the field’s maturity and its continued relevance to pressing questions about how cooperation and conflict shape complex systems. Yet, this rapid, distributed growth has created its own challenges.
@@
This book provides a comprehensive, systematic treatment of signed networks that consolidates foundational theories, clarifies methodological choices, identifies persistent misconceptions, and discusses paths forward. We treat signed networks as a unified analytical framework while respecting that different domains may require different approaches. We emphasise methodological rigour and transparency, discussing assumptions, limitations, and appropriate contexts rather than simply cataloguing methods. We integrate across levels of analysis—static structure and temporal dynamics, processes on and of networks, microscopic mechanisms, and macroscopic patterns. We are selective rather than exhaustive, focusing on foundational concepts, broadly useful methods, and findings that reveal general principles. The book builds systematically. Chapter 2 establishes mathematical foundations: formalising signed networks and presenting the adaptation of classical metrics. Chapter 3 addresses balance theory, distinguishing cognitive from structural balance and presenting measures for quantifying balance. Chapter 4 explores structural inference: null models, embeddings, and algorithms for sign prediction and community detection. Chapter 5 examines dynamics: opinion formation, network evolution, and co-evolutionary frameworks. Chapter 6 tackles data collection and surveys available datasets. Chapter 7 addresses common misconceptions and errors. In Chapter 8, we discuss open challenges and promising directions. Our aim with this book is to make signed network science more accessible without sacrificing rigour, more unified without erasing important distinctions, and more self-aware about its limitations, helping researchers effectively use this powerful lens for understanding how positive and negative relationships jointly shape complex systems.
Figure 1: Examples of signed networks across different domains. In social systems, nodes represent individuals and edges denote friendship/enmity or trust/distrust relationships. In ecological systems, edges capture competitive, mutualistic, or predator–prey dynamics. In economic systems, assets are connected based on correlations between their price time series. In psychological systems, nodes represent beliefs or attitudes, and signed edges encode whether they reinforce or contradict each other. In biological systems, interactions between genes and proteins can be represented through activation/inhibition links. In neuroscience and molecular systems, signed networks can represent either excitatory/inhibitory processes, such as synapses between individual neurons, or functional relationships between brain regions. Finally, in international relations, countries may be connected by alliances or conflicts. Figures adapted from [101, 4, 260, 398, 90].
@@ 2.1 Definition and representation of signed networks
Let us introduce the fundamental mathematical concepts and terms, and some of the key concepts used throughout this book. From a mathematical point of view, networks (whether signed or unsigned) are modelled as graphs [296]. A graph $G=(V,E)$ consists of a set $V=\{v_{i}\}$ of nodes (also called vertices) and a set $E=\{e_{ij}\}$ of edges (also called links) that connect pairs of nodes. An edge connecting the nodes $v_{i}$ and $v_{j}$ is denoted by $e_{ij}=(v_{i},v_{j})$. The number of nodes is denoted by $n$ and the number of edges by $m$. Nodes can represent entities such as people, genes, countries, or more abstract constructs like stocks or time series. Edges capture pairwise interactions between these nodes: social connections between individuals, regulatory relationships between genes, or correlations between time series, among others.
On a signed graph, edges carry a sign attribute [412] expressing a negative or positive type of relation between the nodes. Beyond signs, edges may also have directionality to allow for the modelling of asymmetrical relations between nodes [254], and weights to quantify varying levels of importance across relations [32, 295]. The interplay between these attributes can lead to rich emergent phenomena; for instance, biological circuits leverage both sign and directionality to generate oscillatory dynamics [8]. In this book, we focus on the specific role of signed edges, and we avoid details regarding the role of directionality and weights.
@@
The adjacency matrix enables the application of techniques from linear algebra and spectral graph theory to analyse graphs [38, 413]. When the graph is unsigned, the $(i,j)$th element of $A^{k}$ counts the number of walks of length $k$ between nodes $v_{i}$ and $v_{j}$. When the graph is signed, it counts the difference between the number of positive and negative walks. These are given respectively by
@@ 2.2 Basic metrics
Centrality measures allow us to characterise local and global properties of networks. However, typical network metrics, such as degree, centrality measures, shortest-path distances, clustering coefficients, and network motifs [296], are not straightforward to generalise for signed networks. Some of these cease to be well-defined (e.g., the shortest-path distance), while others admit different generalisations suitable to different contexts. In this section, we explore how to approach each of these metrics and centrality measures in signed networks in a principled manner.
@@ 2.2.1 Signed degree and its distribution
In an unsigned network, the degree of a node $v_{i}$, denoted by $k_{i}$, counts the number of edges adjacent to the node. In a social network, the degree usually quantifies the popularity or influence of a person. Studying the distribution of degrees $P(k)$ can then help answer questions such as whether social influence is concentrated in the hands of a few or more equally spread in the population. When the network is signed, the degree is not characterised by a single number, but by a pair $(k^{+},k^{-})$, denoting the number of positive and negative edges incident to the node, respectively. The two-dimensional nature of the degree causes the degree distribution of a signed network to become bivariate $P(k^{+},k^{-})$. Analysing or visualising bivariate distributions can be challenging, so it is common to consider the marginal distributions, $P(k^{+})$ and $P(k^{-})$. However, one should keep in mind that marginalising erases any correlation between $k^{+}$ and $k^{-}$, which may contain valuable structural information. An alternative is to analyse the absolute or underlying degree $k=k^{+}+k^{-}$, and the net or signed degree $k_{\text{net}}=k^{+}-k^{-}$, through their joint probability distribution $P(k,k_{\text{net}})$ and its marginals. Asymmetries have been found between the distributions $P(k^{+})$ and $P(k^{-})$ [371], suggesting fundamental differences in the formation mechanisms of positive and negative links, and challenging the use of $P(k)$ as an appropriate coarse-graining.
@@ 2.2.2 Signed clustering coefficient
The (unsigned) clustering coefficient measures local cohesiveness by quantifying the abundance of triangles. This is a key structural property of networks that plays a central role in the famous small-world phenomenon [395]. For the signed case, the (global) clustering coefficient can be computed as [223]:
@@
where $k_{i}$ is the absolute degree of node $v_{i}$. This index captures the difference between the number of positive triangles $t^{+}$ and negative triangles $t^{-}$, with the remaining terms [^1] In the presence of self-loops, one must use the normalisation constant $k_{i}^{2}/2$ instead [223]. ensuring that it remains bounded between +1+1 and 1-1. Clustering can also be studied at the node level, quantifying the difference between the number of positive triangles ($t^{+}$) and negative triangles ($t^{-}$) containing node $v_{i}$, relative to all two-paths passing through $v_{i}$ [76]:
@@
Both $C$ and $C_{i}$ penalise negative triangles ($t^{-}$) and unclosed triangles. To isolate the effect of negative triangles while controlling for network topology, Kunegis introduced the relative clustering coefficient [223], defined as the ratio between the signed and unsigned clustering coefficients:
@@ 2.2.3 Signed distances
Another fundamental concept in network analysis is the distance between pairs of nodes. This measure plays a central role in various applications, including the definition of various centrality metrics [296, 47], community detection [293], and network geometry [43]. Formally, a distance is a function that takes two nodes as input and returns a real number that satisfies four key properties.
@@
In the case of signed networks, however, these measures fail to satisfy the properties of a distance [93]. To understand the reasons, consider a graph made of three nodes $v_{1},v_{2},v_{3}$, the first two linked by a negative edge and the last two by a positive one. Using the minimum-weight distance we obtain $d(v_{1},v_{2})<0$, violating the non-negativity condition, and $d(v_{1},v_{3})=0$, contradicting the identity of indiscernibles. These issues are not just theoretical; they can result in algorithms selecting infinitely long paths as optimal routes.
A natural way to overcome these limitations is to use generalised distances [43]. These distances extend beyond simple shortest-path measures by incorporating the collective influence of all paths between two nodes, often through dynamic processes such as diffusion, random walks, or epidemic spreading. One prominent example is the resistance distance, which originated from electrical circuit theory [61]. Here, the distance between two nodes corresponds to the effective resistance in a hypothetical circuit where a unit current (1 amp) is injected into $v_{i}$ and extracted from $v_{j}$, with every edge acting as a 1-ohm resistor. Alternatively, this measure has a probabilistic interpretation: it is equal to the expected commuting time for a random walker travelling from $v_{i}$ to $v_{j}$ and back. Mathematically, the resistance distance is given by
@@
where $L$ is the graph Laplacian [87], discussed in detail in Section 2.3, and $L^{\dagger}$ is its Moore-Penrose pseudoinverse, defined as: $L^{\dagger}=(L+\gamma J)^{-1}-\gamma^{-1}J$, with $\gamma$ a regularization parameter and $J$ an all-ones matrix [^3] The Moore–Penrose pseudoinverse generalizes the notion of the matrix inverse and remains well defined even when some eigenvalues are equal to zero, as in the Laplacian. For nonsingular matrices, it coincides with the ordinary inverse.. The function $d_{res}$ satisfies all properties of a distance, even when including negative edges [225]. Moreover, it is Euclidean, meaning it can be interpreted as the usual “straight-line length” in some abstract Euclidean space.
The effective resistance is not the only generalised distance that retains good properties in signed networks. Other relevant ones include the diffusion distance [80], which captures how similar two nodes are based on the spread of a diffusive process over time; and the communicability distance [108, 92], which measures the ease of information transfer between nodes using the exponential of the adjacency matrix.
Despite the mathematical challenges discussed, algorithms exist to compute the minimum-weight distance in signed directed graphs. Two important examples are the Bellman-Ford and Floyd-Warshall algorithms [75]. Both algorithms are based on dynamic programming and are preferred over Dijkstra’s algorithm when negative edges are present. Apart from computing minimum-weight paths, these algorithms also detect $\Sigma$-negative cycles (cycles where the sum of edge weights is negative) [^4] Note that this notion is different from negative cycles, where the product of edge weights is negative. $\Sigma$-negative cycles appear in diverse applications. For instance, in exchange-rate networks, currencies can be represented as nodes and exchange rates as edge weights (typically using the negative logarithm of the rate). In this setting, a $\Sigma$-negative cycle indicates an arbitrage opportunity..